Capicator

Capicator

Grade:12

3 Answers

askiitian expert akshay singh
14 years ago

hi aman,

   i think the ques is incomplete so plz send complete ques and then let me help u

Aman Deep Singh
14 years ago
My mean to say is that if we connect a capacitor in between number of resistors than what will be effect on its net resistance
Pratham Ashish
14 years ago

hi aman,

when u connect capacitor wid resistors den then the capacitor would discharge its energy into the resistor. This voltage across the capacitor over time could be found through Kirchhoff's current law, where the current coming out of the capacitor must equal the current going through the resistor. This results in the linear differential equation

C\frac{dV}{dt} + \frac{V}{R}=0.

Regarding its equivalent resistance,

the equivalent resistance of a capacitor increases in relation to the amount of charge stored on the capacitor. If a capacitor is subjected to an alternating current voltage source, then the voltage of the capacitor would flip to the frequency of the AC voltage source. The faster the voltage of the AC voltage source flips, the less time charge would be allowed to be stored on the capacitor, therefore reducing the capacitor's equivalent resistance. This explains the inverse relationship the equivalent resistance of a capacitor has with the frequency of the voltage source.

The resistance, also known as the complex impedance, ZC (in ohms) of a capacitor with capacitance C (in farads) is

s is d angular frequency

Z_C = \frac{1}{sC}

The current in the circuit is the same everywhere since the circuit is in series:

I(s) = \frac{V_{in}(s) }{R+1/ Cs}  =  { Cs \over 1 + RCs } V_{in}(s)

u can see dat in d expression of current resistance of capacitor just adds up to d given resistance...